# TEn-CATS: Text-Enriched Audio-Visual Video Parsing with Multi-Scale Category-Aware Temporal Graph

### v1

@@ IV-C1 Ablation Study on Module Contribution
To verify the contribution of the proposed BiT and CATS modules, we conduct ablation studies by disabling these two components in TEn-CATS. To ensure a fair comparison with CoLeaF, which was originally built upon audio and visual features extracted by VGGish and ResNet, we re-train CoLeaF using the same input features as TEn-CATS, denoted as CoLeaF.
To validate the effectiveness and individual contribution of each proposed module, we first conduct ablation studies on the LLP, a benchmark dataset for AVVP task. Specifically, we compare the performance of TEn-CATS when removing or isolating the BiT and CATS modules, as shown in Table II.
Overall, we observe that the best results for A and Event@AV at both segment and event levels are achieved when only the BiT module is applied, while the best results for V, AV, and Type@AV at both levels are obtained when combining both BiT and CATS (full model). This reveals a clear trade-off of our model between enhancing unimodal performance, especially audio-driven event recognition, and maximizing overall multi-modal and event consistency. For the best performance in A and Event@AV when only adding BiT module, this is probably because the BiT-only configuration focuses on strengthening the semantic alignment between modality-specific features and event concepts without introducing additional temporal constraints. This simplifies the learning objective, allowing the model to better capture highly distinguishable audio characteristics, such as speech patterns, musical rhythms, environmental sounds (e.g., alarm, dog barking), or other acoustically unique event cues. These audio events are often globally consistent across segments, making them easier to capture through semantic enhancement alone, without requiring complex temporal reasoning. Moreover, this semantic reinforcement directly benefits holistic event detection (Event@AV) by helping the model capture high-level event semantics across segments, even if temporal precision is not fully optimized.
Focusing on the performance of our framework, the full model introduces category-aware temporal modeling, which is particularly effective for visual streams, where events often happen with clear temporal structures (e.g., object actions, scene changes). The combination of semantic and temporal cues also benefits AV fusion and type-level performance (Type@AV), as it helps the model align and balance both modalities over time. However, this more complex modeling may slightly dilute unimodal audio strength, as the model focuses more on multi-modal consistency rather than maximizing single-modality performance.
TABLE II: Ablation study for TEn-CATS. o. means only have certain module, A and V means audio branch and visual branch.
@@
CATS o. A 65.5 68.1 59.6 64.4 64.5
CATS o. V 66.0 68.9 60.4 65.1 65.2
CATS full 65.3 68.7 60.8 64.9 64.6
@@
CATS o. A 54.5 64.8 51.5 56.9 54.0
CATS o. V 55.1 65.4 52.2 57.6 54.7
CATS full 54.7 64.8 52.8 57.5 54.3
@@ IV-C2 Qualitative Case Study on the BiT Module
To qualitatively assess the impact of the BiT module, we visualize a representative example from the LLP dataset in Fig. 6. This case illustrates the predicted temporal segments of the audio and visual events by different models, including CMPAE [18], CoLeaF [19], and our model with only the BiT module enabled.
@@
IV-C3 Analysis of Temporal Scale in the CATS Module
To evaluate the sensitivity of our CATS module to the choice of the temporal hop size $k$, we conduct controlled experiments by varying $k\in{2,3,4,5}$. Fig. 7 presents the performance across five key metrics (A, V, AV, Type@AV, Event@AV) at both the segment-level and event-level, color-coded for each metric group.
@@ IV-C3 Analysis of Temporal Scale in the CATS Module
Figure 7: Segment- and event-level performance of the CATS module under different hop sizes $k$. Each color represents one evaluation metric (A, V, AV, Type@AV, Event@AV).
TABLE III: Comparison of CATS performance on the weakly-labeled UnAV-100 dataset.
@@
CATS 41.9 (+0.4) 47.5
IV-C4 Effectiveness of CATS on UnAV-100
To evaluate the generalizability and robustness of our proposed CATS module, we conduct additional experiments on the UnAV-100 dataset using the weak labels provided by CoLeaF. As shown in Table III, the incorporation of CATS yields further improvement over the previous state-of-the-art methods.
Compared to CoLeaF, our method achieves 41.9% segment-level mAP and 47.5% event-level mAP in the AV branch, surpassing it by +0.4% and maintaining competitive performance on the event level. This gain, albeit marginal, is consistent and meaningful under weak supervision, demonstrating that CATS can effectively model category-aware temporal structures even when ground-truth timestamps are unavailable.
In particular, the improvement on the segment-level metric highlights CATS’s ability to enhance temporal localization and reduce false positives in ambiguous segments. These results further validate that our framework generalizes well across different weakly labeled datasets and benefits from the fine-grained temporal reasoning introduced by the CATS module.
@@ V Conclusion
We have presented a novel framework for AVVP task by designing the BiT and CATS modules. Extensive experiments on LLP and UnAV-100 demonstrate the effectiveness of our approach. BiT enhances audio-visual understanding, especially for ambiguous or rare events, while CATS ensures consistent semantic reasoning over time. Ablation studies further validate the individual contributions and the strong complementarity of both modules. Our model achieves new state-of-the-art results and shows robust generalization under real-world conditions. Future work includes handling overlapping events, enhancing robustness of the models to label noise, and extending to tasks such as audio-visual instance segmentation and cross-modal retrieval.

### v2

@@ IV-C1 Ablation Study on Module Contribution
To verify the contribution of the proposed BiT and CATG modules, we conduct ablation studies by disabling these two components in TEn-CATG. To ensure a fair comparison with CoLeaF, which was originally built upon audio and visual features extracted by VGGish and ResNet, we re-train CoLeaF using the same input features as TEn-CATG, denoted as CoLeaF. To validate the effectiveness and individual contribution of each proposed module, we first conduct ablation studies on the LLP dataset. Specifically, we compare the performance of TEn-CATG when removing or isolating the BiT and CATG modules, as shown in Table II.
Overall, we observe that the best results for A and Event@AV at both segment and event levels are achieved when only the BiT module is applied, while the best results for V, AV, and Type@AV at both levels are obtained when combining both BiT and CATG (full model). This reveals a clear trade-off of our model between enhancing unimodal performance, especially audio-driven event recognition, and maximizing overall multi-modal and event consistency. For the best performance in A and Event@AV when only adding the BiT module, this is probably because the BiT-only configuration focuses on strengthening the semantic alignment between modality-specific features and event concepts without introducing additional temporal constraints. This simplifies the learning objective, allowing the model to better capture highly distinguishable audio characteristics, such as speech patterns, musical rhythms, environmental sounds (e.g., alarm, dog barking), or other acoustically unique event cues. These audio events are often globally consistent across segments, making them easier to capture through semantic enhancement alone, without requiring complex temporal reasoning. Moreover, this semantic reinforcement directly benefits holistic event detection (Event@AV) by helping the model capture high-level event semantics across segments, even if temporal precision is not fully optimized.
Focusing on the performance of our framework, the full model introduces category-aware temporal modeling, which is particularly effective for visual streams, where events often happen with clear temporal structures (e.g., object actions, scene changes). The combination of semantic and temporal cues also benefits AV fusion and type-level performance (Type@AV), as it helps the model to align and balance both modalities over time. However, this more complex modeling may slightly dilute unimodal audio strength, as the model focuses more on multi-modal consistency rather than maximizing single-modality performance.
TABLE II: Ablation study for TEn-CATG. o. means only have certain module, A and V means audio branch and visual branch.
@@
CATG o. A 65.5 68.1 59.6 64.4 64.5
CATG o. V 66.0 68.9 60.4 65.1 65.2
CATG full 65.3 68.7 60.8 64.9 64.6
@@
CATG o. A 54.5 64.8 51.5 56.9 54.0
CATG o. V 55.1 65.4 52.2 57.6 54.7
CATG full 54.7 64.8 52.8 57.5 54.3
@@ IV-C2 Qualitative Case Study on the BiT Module
To assess the impact of the BiT module, we visualize a representative example from the LLP dataset in Fig. 6. This case illustrates the predicted temporal segments of the audio and visual events by different models, including CMPAE [18], CoLeaF [19], and our model with only the BiT module enabled.
@@
IV-C3 Analysis of Temporal Scale in the CATG Module
To evaluate the sensitivity of our CATG module to the choice of the temporal hop size $k$, we conduct controlled experiments by varying $k\in{2,3,4,5}$. Fig. 7 presents the performance across five key metrics (A, V, AV, Type@AV, Event@AV) at both the segment-level and event-level, color-coded for each metric group.
@@ IV-C3 Analysis of Temporal Scale in the CATG Module
Figure 7: Segment- and event-level performance of the CATG module under different hop sizes $k$. Each color represents one evaluation metric (A, V, AV, Type@AV, Event@AV).
TABLE III: Comparison of CATG performance on the weakly-labeled UnAV-100 dataset.
@@
CATG 41.9 (+0.4) 47.5
IV-C4 Effectiveness of CATG on UnAV-100
To evaluate the generalizability and robustness of our proposed CATG module, we conduct additional experiments on the UnAV-100 dataset using the weak labels provided by CoLeaF. As shown in Table III, the incorporation of CATG yields further improvement over the previous state-of-the-art methods.
Compared to CoLeaF, our method achieves 41.9% segment-level mAP and 47.5% event-level mAP in the AV branch, surpassing it by +0.4% and maintaining competitive performance on the event level. This gain, albeit marginal, is consistent and meaningful under weak supervision, demonstrating that CATG can effectively model category-aware temporal structures even when ground-truth timestamps are unavailable.
In particular, the improvement on the segment-level metric highlights CATG’s ability to enhance temporal localization and reduce false positives in ambiguous segments. These results further validate that our framework generalizes well across different weakly labeled datasets and benefits from the fine-grained temporal reasoning introduced by the CATG module.
@@ V Conclusion
We have presented a novel framework for AVVP task by designing the BiT and CATG modules. Extensive experiments on LLP and UnAV-100 datasets demonstrate the effectiveness of our approach. BiT enhances audio-visual understanding, especially for ambiguous or rare events, while CATG ensures consistent semantic reasoning over time. Ablation studies further validate the individual contributions and the strong complementarity of both modules. Our model achieves new state-of-the-art results and shows robust generalization under real-world conditions. Future work includes handling overlapping events, enhancing robustness of the models to label noise, and extending to tasks such as audio-visual instance segmentation and cross-modal retrieval.